A 0.5Kg rider travelling to the right at 2ms-1 collides with a stationary rider of mass 1Kg. Find the velocities of the two riders after the collision if: a. It is perfectly elastic.

1
Expert's answer
2021-10-26T09:38:26-0400

"m_1 = 0.5kg; v_1=2\frac{m}{s}"

"m_2 = 1kg;v_2=0"

"\text{Since a perfectly elastic collision:}" "\text{}"

"m_1v_1+m_2v_2 = m_1u_1+m_2u_2\ (1)"

"\frac{m_1v_1^2}{2}+\frac{m_2v_2^2}{2}=\frac{m_1u_1^2}{2}+\frac{m_2u_2^2}{2}\ (2)"

"\text{from }(1)"

"0.5*2+1*0= 0.5u_1+1*u_2"

"u_2= 1-0.5u_1\ (3)"

"\text{from }(2)"

"\frac{0.5*2^2}{2}+\frac{1*0^2}{2}=\frac{0.5*u_1^2}{2}+\frac{1*u_2^2}{2}"

"u_2^2 = 2-0.5u_1^2\ (4)"

"\text{from (3) and (4)}"

"u_2^2>0"

"2-0.5u_1^2>0"

"u_1^2<4"

"(1-0.5u_1)^2=2-0.5u_1^2"

"0.75u_1^2-u_1-1 =0"

"u_1\approx -0.67\frac{m}{s}"

"u_1'=2(\text{does not satisfy the condition }u_1^2<4)"

"u_2= 1-0.5u_1= 1+0.5*0.67\approx 1.34\frac{m}{s}"


"\text{Answer:}"

"-0.67\frac{m}{s}\text{ first rider speed after collision}"

"1.34\frac{m}{s}\text{ second rider speed after collision}"